OpenStudy (anonymous):

simplify the expression (cos4x + cos3x) / (sin4x - sin3x)

OpenStudy (jdoe0001):

http://www.sosmath.com/trig/Trig5/trig5/img9.gif \(\bf \cfrac{cos(4x)+cos(3x)}{sin(4x)-sin(3x)}\\ \quad \\ \quad \\ \cfrac{ \cancel{2cos\left(\frac{4x+3x}{2}\right)}cos\left(\frac{4x-3x}{2}\right) }{ \cancel{2cos\left(\frac{4x+3x}{2}\right)}sin\left(\frac{4x-3x}{2}\right) } \implies \cfrac{cos\left(\frac{4x-3x}{2}\right)}{sin\left(\frac{4x-3x}{2}\right)}\implies tan\left(\frac{4x-3x}{2}\right)\\ \quad \\ tan\left(\frac{x}{2}\right)\)

OpenStudy (jdoe0001):

hmmm actually

OpenStudy (jdoe0001):

\(\bf \cfrac{cos(4x)+cos(3x)}{sin(4x)-sin(3x)}\\ \quad \\ \quad \\ \cfrac{ \cancel{2cos\left(\frac{4x+3x}{2}\right)}cos\left(\frac{4x-3x}{2}\right) }{ \cancel{2cos\left(\frac{4x+3x}{2}\right)}sin\left(\frac{4x-3x}{2}\right) } \implies \cfrac{cos\left(\frac{4x-3x}{2}\right)}{sin\left(\frac{4x-3x}{2}\right)}\implies cot\left(\frac{4x-3x}{2}\right)\\ \quad \\ cot\left(\frac{x}{2}\right)\)

OpenStudy (anonymous):

THANK YOU SOOOOOO MUCH!!!!!!!!!!! THAT IS AMAZING!!!! :D

OpenStudy (anonymous):

how did you get 2 cos (4x+3x)/2?

OpenStudy (jdoe0001):

http://www.sosmath.com/trig/Trig5/trig5/img9.gif <---- from the sum to product trigonometry identities